Hooker Company issues $200,000 of ten-year, 8% bonds to yield 10% on January 1, 2018. The bonds pay interest annually on December 31. The bonds were sold at a discount of $24,578.

The bond interest expense for 2018 is:

Multiple Choice

A) $16,000.

B) $17,542.

C) $20,000.

D) $21,542.
Answer #1

Answers : B) $ 17,542

Interest Expenses = Bond Carrying Value at the beginning of the Period * Market Interest Rate

Bond Carrying Value = 200,000-24,578 =175,422

Market interest rate = 10%

For 2018

Bond Interest Expenses = 175,422 *10%-= $17,542 (Answer)